OpenStudy (anonymous):

When two gears are mashed, the revolutions per minute (rpm) are inversely proportional to the number of teeth, as shown below. In one machine the rpm ratio of the large gear to the small gear is 1:3. the small gear has 44 teeth. In a second machine the large gear revolves at 200 rpm and has 423 teeth. The small gear revolves at 900 rpm rpm of large gear/rpm of small gear= # of teeth on small gear/ teeth on large gear

OpenStudy (ranga):

There is no question here.

OpenStudy (anonymous):

I need help finding the number of teeth on the large gear of the first machine and then also finding the number of teeth on the small gear of the second machine

OpenStudy (anonymous):

I have no idea how to even begin to find the answer

OpenStudy (ranga):

In one machine the RPM ratio of the LARGE gear TO the SMALL gear is 1:3. rpm_large / rpm_small = 1 / 3 Since rpm is inversely proportional to the number of teeth: rpm_large / rpm_small = teeth_small / teeth_large ----- (1) The small gear has 44 teeth. Plug in the numbers into (1): 1 / 3 = 44 / teeth_large cross multiply: teeth_large = 44 x 3 = 132 teeth on the larger gear.

OpenStudy (ranga):

You want to try the second part?

OpenStudy (anonymous):

200/900?

OpenStudy (ranga):

That is the ratio on the left. You have correctly found rpm_large / rpm_small = 200 / 900 Now do the right hand side.

OpenStudy (anonymous):

so the first machine has 132 teeth on the large gear?

OpenStudy (ranga):

Yes, the first machine has 132 teeth on the large gear.

OpenStudy (anonymous):

3.5/423?

OpenStudy (ranga):

rpm_large = 200 rpm_small = 900 teeth_large = 423 teeth_small = ? Plug into the formula: rpm_large / rpm_small = teeth_small / teeth_large 200 / 900 = teeth_small / 423 multiply both sides by 423 200 * 423 / 900 = teeth_small teeth_small = 94

OpenStudy (anonymous):

that makes more sense how you wrote it out. I was confused

OpenStudy (ranga):

alright.

OpenStudy (anonymous):

so you did 200*423 which = 84600 and you took that and divided it by 900? and that is how you got 94?

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

ok thank you so much

OpenStudy (ranga):

you are welcome.